Alan and Jim were asked by their boss to work on a project. Since Jim had to visit a company's factory outside of the state, Alan started the project alone. After 10 days Jim came back and they worked for 5 days to finish the project. If Alan could have finished the project in 23 days when Alan worked alone from the beginning to the end, how long would it take if Jim had to work alone from the bedid to the end to finish the project? Your answer is supposed to be d days and h hours where d is a positive integer and h is a positive integer or 0.
Answer by mukhopadhyay(490)

Alan can finish the same job in 23 days
=> Alan completes 1/23 portion of the job in 1 day
Alan worked by himself for 10 days and he aslo spent an additional 5 days with Jim to finish the job
So, Alan spent a total of 15 days for the project
In 15 days, Alan did 15/23 portion of the job
The remaining 8/23 (1 - 15/23) portion of the job was done by Jim in 5 days
Jim does 8/23 portion of the job is 5 days
=> Jim does 1 job is 5*(23/8) = 115/8 days
115/8 days = 14 and 3/8 days = 14 days and 24(3/8) hours = 14 days and 9 hours
So, if Jim had to finish the project from start to end all by himself, it would have taken him 14 days 9 hours to do so.